# Uncertainty of Thoughts: Uncertainty-Aware Planning Enhances Information Seeking in Large Language Models

### v2

@@ 1 Introduction
As the capabilities of large language models (LLMs) grow, they are being increasingly deployed in challenging real-world settings involving uncertainty and ambiguity. In particular, recent work aims to develop LLM agents or assistants xi2023rise ; park2023generative that effectively complete tasks in interactive environments, leading to a growing need for LLMs that can actively seek the information they need to solve a task by asking questions in conversational settings. For example, in medical diagnosis, patients often do not initially report their symptoms in full detail. In such situations, the ability of the doctor to ask effective questions is crucial, as a successful diagnosis often depends on revealing important details that the patient did not initially provide (Figure 1).
@@
Recent techniques aim to improve LLMs’ reasoning or planning abilities based on the given information rather than enabling LLMs to seek information efficiently. For example, Chain-of-Thought (CoT) (wei2022chain, ) and Tree-of-Thoughts (ToT) (yao2023tree, ) allow LLMs to express intermediate ‘thoughts’ and reason over them. Unlike these methods, our focus is on enabling the LLM to ask questions effectively by explicitly guiding the model toward reducing uncertainty, which these do not consider. Thus, they lack effective signals for what questions can better reduce model uncertainty by uncovering critical information.
@@
Additionally, most standard benchmarks for LLMs, particularly in question answering, assume that all necessary information to solve a task is provided at the outset, and thus do not evaluate the model’s active information-seeking capabilities. To close this gap, we first introduce a benchmark comprising 5 datasets [^2] We also incorporate the efforts of prior datasets srivastava2022beyond ; xu2019end ; liu2022meddg ; raghu-etal-2021-end , through further work and refinement to construct this benchmark. Details are introduced in section 3 Experiments and Appendix I.2. on 3 tasks: 20 Questions, a simplified medical diagnosis task, and a basic troubleshooting task. These tasks are designed to measure the model’s ability to ask questions effectively to gather the information they need. For example, the 20 Questions game, also studied by Noever et al.noever2023chatbots , requires the model to ask ‘yes’ or ‘no’ questions to determine an unknown object or entity. This scenario serves as a clear and easily analyzed test case, isolating the model’s ability to recognize its own uncertainty, and to ask questions that guide it to the correct answer.
@@ 2.4 Uncertainty-Based Reward Calculation
Entropy. Entropy and information gain are well-known concepts in information theory shannon1948mathematical . In our work, we use these concepts to measure how much information is gained (or equivalently, how much uncertainty is reduced) by asking a question, to formulate our rewards. Entropy measures the level of uncertainty in a random variable: higher entropy indicates greater uncertainty. The entropy of a discrete random variable $X$ taking values $x_{1},...,x_{n}$ is:
@@
$H_{v}(X):=\sum\nolimits_{i:\omega_{i}\in\Omega_{v}}p(x_{i}|\Omega_{v})\log p(x_{i}|\Omega_{v})\vspace{-4mm}$ (5)
@@
$H_{v}^{A}(X):=\sum\nolimits_{i:\omega_{i}\in\Omega_{v}^{A}}p(x_{i}|\Omega_{v}^{A})\log p(x_{i}|\Omega_{v}^{A})\vspace{-2mm}$ (6)
@@
$\vspace{-3mm}R_{u}(v)=\widetilde{IG}_{v}(X):=(-p_{v}^{A}\log p_{v}^{A}-p_{v}^{N}\log p_{v}^{N})/(1+\lambda^{-1}|p_{v}^{A}-p_{v}^{N}|)\vspace{-3mm}$ (10)
@@ 2.5 Question Selection Via Reward Propagation
$\vspace{-3mm}R_{e}(v):=$$\begin{cases}R_{a}(v)&\text{if $v$ is a leaf; otherwise:}\\ p_{v}^{A}R_{e}(v^{A})+p_{v}^{N}R_{e}(v^{N})&\text{if $v$ is an Answerer Node}\\ \frac{1}{m}\sum_{w\in\mathsf{Children}(v)}^{m}R_{e}(w)&\text{if $v$ is a Questioner Node}\\ \end{cases}\vspace{-3mm}$
@@ 2.6 UoT Summary
UoT generates candidate questions $q_{i}^{1},q_{i}^{2},\dots,q_{i}^{m}$ based on history and the current possibility set $\Omega_{i}$. It simulates a tree for each question, calculates uncertainty-based rewards $R_{u}(v)$, and computes expected rewards $R_{e}(v)$. The question $q_{i}^{n}$ with the highest expected reward is chosen for interaction.
@@ 2.7 Extensions and Discussion
Open Set UoT. Recall that in the closed set scenario, the Questioner starts with knowledge of the possibility space $\Omega$. In practice, the possibility space is often unknown, resulting in the open set setting. To adapt UoT to this case, we prompt Questioner to initialize the possibility space $\Omega$ and then reinitialize the possibility set $\Omega_{i}$ according to current history $h_{i}$. Then, the rest of UoT is unchanged. The generalization in open-end answers. The UoT framework enables LLMs to update possibilities after each interaction, including affirmative/negative or open-ended responses. Thus, it can be applied to open-ended answers scenarios. Pruned UoT. To enhance simulation efficiency, pruning similar to Beam Search can be used to limit the number of explored paths in the simulation trees to a fixed size.
@@ 3.1 Experimental Setup
We test various LLMs to evaluate the generality of our method, including Llama-3-70B-Instruct llama3modelcard , Mistral-Large mistrallarge , Gemini-1.5-Pro reid2024gemini , Claude-3-Opus claude3 and GPT-4 openai2023gpt4 . We also validate the performance of earlier released LLMs (Refer to Appendix D) including Llama 2-70B-Chat touvron2023llama , Cohere cohere , PaLM 2 anil2023palm , Claude 2 Claude-2 and GPT-3.5-turbo openai2023gpt3.5 .
BaselinesDirect Prompting (DP) prompts an LLM directly to generate the next response. Planning Prompting (PP) is motivated by Wang et al.wang2023plan . We leverage another LLM to plan the future and, consequently, determine the question to ask. Chain-of-Thought (CoT) wei2022chain improves reasoning in LLMs by detailing reasoning steps. CoT-SC (Self-Consistency) wang2022self samples various paths to enhance fairness in computational comparisons. Reflexion shinn2023reflexion lets agents propose actions and self-assess to foster new ideas. Tree-of-Thoughts (ToT) yao2023tree enables LLMs to make decisions by exploring and evaluating multiple reasoning paths over a tree structure. We examine ToT under two setups: Original-ToT, which uses the standard approach of generating and evaluating questions, and Adapted-ToT (Ad.-ToT), where we integrate heuristic experience into prompt for question generation and evaluation, focusing on questions that halve the search space. We matched the tree depth to the simulation steps in our UoT method for a fair comparison. We evaluate methods and LLMs in both open set (OS) and closed set (CS) settings. In open set, models are tested without prior knowledge of outcomes; in closed set, they are given complete information about all possible outcomes. For details, see Appendix I.1 for experimental settings and Appendix L for prompts.
@@
Scenarios and Datasets20 Questions is a game where the answerer thinks of an item and the questioner asks up to 20 yes-or-no questions to guess it. We use two datasets, Common (collected by us, refer to Appendix 6 for more details) and Things hebart2019things , including 111 and 1854 items separately. In this scenario, the maximal turns is set to 20. In Medical Diagnosis, the doctor needs to ask questions to patients about their symptoms, to determine an accurate diagnosis. We use two datasets: DX xu2019end , with 104 doctor-patient dialogues and 5 diseases in test set, and MedDG liu2022meddg with over 17K conversations across 15 disease types. We manually selected 500 high-quality samples for evaluation (see Appendix I.3 for selection process). Importantly, Open-ended responses from patient are allowed in MedDG to validate UoT’s generalization in open-ended scenarios. Both datasets are limited to 5 turns. Troubleshooting involves a technician working with customers to identify and fix problems in computer systems, electronic devices, or machinery. Raghu et al.raghu-etal-2021-end introduce FloDial with 894 dialogues, containing 153 faults. We evaluate using a maximum of 20 turns. The answerer, simulated by GPT-4, is prompted with the patient’s actual disease and conversation details for each case. For more details, refer to Appendix I.2 and see examples of these scenarios in Appendix K.
UoT (Open Set) Setup We iteratively update LLMs’ perceived possibilities based on conversational history, rather than defining them all upfront. In medical diagnosis and troubleshooting, initial descriptions from symptoms or issues help set up initial possibilities. In the 20-question game, we start with broad inquiries using the Direct Prompting method for the first three rounds to gather more information. The ToT tree structure method employs a similar strategy. Setup details in Appendix I.4.
Evaluation Metrics  To measure efficacy and efficiency, we use: Success Rate (%): $\textbf{SR}=S/T$, where $S$ is the number of successful cases, and $T$ is the total number of cases; Mean Conversation Length in Successful Cases: $\textbf{MSC}=R_{s}/S$, where $R_{s}$ is the total rounds in successful cases; Mean Conversation Length: $\textbf{MCL}=R/T$, where $R$ is the total rounds in all cases.
@@ 3.2 Performance
20 Questions  As illustrated in Table 5, for all types of LLMs, those equipped with UoT outperform the baselines in both open set and close settings. Among the methods used on GPT-4 to enhance planning and reasoning, CoT (CS) and PP (CS) show inferior performance even compared to GPT-4 alone. UoT (OS) demonstrates superior performance, with with an average 8.7% improvement than Adapted-ToT (OS) in success rate. Moreover, UoT (CS) achieves the highest success rate, surpassing the second-best Reflexion by an average of 4.3%.
Medical Diagnosis  UoT (CS) outperforms baselines in simplified medical diagnostics, achieving a 97.0% success rate on the DX dataset with GPT-4. On the MedDG dataset, UoT (CS) on Gemini-1.5-Pro and GPT-4 achieve success rates of 81.4% and 88.0%. It also reduces conversation lengths to an average MSC of 2.0 on GPT-4 for DX, lower than 3.5 and 3.0 for DP methods. These results demonstrate the versatility of our UoT in handling both binary and open-ended interactions effectively.
Troubleshooting  UoT (CS) with GPT-4 similarly achieves the highest SR of 67.3%, and the lowest MSC of 7.8. It also shows a remarkable improvement from 43.7% to 67.3% in Success Rate.
Overall Performance  On average, UoT enhances the success rate by 38.1% compared to DP across 5 datasets and 5 different LLMs, including open source and commercial models. Notably, Success Rate increases 46.6% for Llama3-70B. Furthermore, UoT outperforms CoT-SC by 33.8% and Reflexion by 29.9%. Even compared to tree structure methods like Original-ToT and Adapted-ToT, UoT still shows superior performance with gains of 28.3% and 12.4% respectively. Additionally, Pruned UoT, our pruning method to improve efficiency, outperforms Adapted-ToT by 7.36%. Additionally, our study shows that UoT’s one-step planning is effective due to effective reward design and question selection. We limit simulations to three steps for budgetary reasons, balancing efficiency and effectiveness. For more details, see Appendix E.
@@
Case Studies and Reliability of GPT-4 as answerer  Figure 3 shows UoT, compared to direct prompting, more effectively reduce uncertainty and narrow down candidates, avoiding overly specific queries. After gaining initial information (e.g., stomach pain), it generates targeted questions about related issues rather than general inquiries. Additionally, GPT-4’s accuracy as answerer is evaluated by analyzing 10% of interactions from each dataset, consistently showing reliable responses. For quantitative details, see Appendix F.
@@ 4 Related Work
Planning and Reasoning of LLMs LLMs show prowess in planning and reasoning. Wei et al.wei2022chain introduced CoT prompting for intermediate reasoning; Yao et al.yao2023tree proposed ToT prompting using DFS/BFS. Besta et al.besta2023graph present GoT to solve elaborate problems. Feng et al.feng2023alphazero illustrated TS-LLM’s tree-search guided decoding. ReAct yao2022react offers acting-based prompting, while Reflexion shinn2023reflexion enhances this with feedback reflection. Zhou et al.zhou2023language unify reasoning and planning.
Decision-making and Information-seeking by LLMs LLMs have evolved as decision-making tools, with models like LLM+P liu2023llm+ and LLM-DP dagan2023dynamic combining external planners and LLMs for natural language-based programming. RAP hao2023reasoning goes beyond structured language, using LLMs with Monte Carlo Tree Search (MCTS) chaslot2008monte for dynamic decision-making. This approach is also seen in the work of Zhao et al.zhao2023large , applying MCTS and LLM knowledge for complex tasks like robot control. However, MCTS struggles in uncertain scenarios due to its reliance on terminal states and specific modules for rewards and action selection. Additionally, to enhance LLMs’ questioning abilities, Deng et al.deng2023rephrase introduce the Rephrase and Respond method. AVIS hu2023avis represents an autonomous visual question answering system that uses external tools. Pan et al.pan2023kwaiagents introduce KwaiAgents for processing queries, following guidelines, and accessing external documents.

### v3

@@ 1 Introduction
As the capabilities of large language models (LLMs) grow, they are being increasingly deployed in challenging real-world settings involving uncertainty and ambiguity. In particular, recent work aims to develop LLM agents or assistants Xi et al. (2023); Park et al. (2023) that effectively complete tasks in interactive environments, leading to a growing need for LLMs that can actively seek the information they need to solve a task by asking questions in conversational settings. For example, in medical diagnosis, patients often do not initially report their symptoms in full detail. In such situations, a doctor’s ability to ask effective questions is crucial, as a successful diagnosis often depends on revealing important details that the patient did not initially provide (Figure 1).
@@
Recent techniques aim to improve LLMs’ reasoning or planning abilities based on the given information rather than enabling LLMs to seek information efficiently. For example, Chain-of-Thought (CoT) (Wei et al., 2022) and Tree-of-Thoughts (ToT) (Yao et al., 2023) allow LLMs to express intermediate ‘thoughts’ and reason over them. Unlike these methods, our focus is on enabling the LLM to ask questions effectively by explicitly guiding the model toward reducing uncertainty, which these do not consider. Thus, they lack effective signals for questions that better reduce uncertainty by revealing critical information.
@@
Additionally, most standard benchmarks for LLMs, particularly in question answering, assume that all necessary information to solve a task is provided at the outset, and thus do not evaluate the model’s active information-seeking capabilities. To close this gap, we first introduce a benchmark comprising 5 datasets [^2] We also incorporate the efforts of prior datasets Srivastava et al. (2022); Xu et al. (2019); Liu et al. (2022); Raghu et al. (2021), through further work and refinement to construct this benchmark. Details are introduced in section 3 Experiments and Appendix I.2. on 3 tasks: 20 Questions, a simplified medical diagnosis task, and a basic troubleshooting task. These tasks are designed to measure the model’s ability to ask questions effectively to gather the information they need. For example, the 20 Questions game, also studied by Noever et al.Noever & McKee (2023), requires the model to ask ‘yes’ or ‘no’ questions to determine an unknown object or entity. This scenario serves as a clear and easily analyzed test case, isolating the model’s ability to recognize its own uncertainty, and to ask questions that guide it to the correct answer.
@@ 2.4 Uncertainty-Based Reward Calculation
Entropy. Entropy and information gain are well-known concepts in information theory Shannon (1948). In our work, we use these concepts to measure how much information is gained (or equivalently, how much uncertainty is reduced) by asking a question, to formulate our rewards. Entropy measures the level of uncertainty in a random variable: higher entropy indicates greater uncertainty. The entropy of a discrete random variable $X$ taking values $x_{1},...,x_{n}$ is:
@@
$H_{v}(X):=\sum\nolimits_{i:\omega_{i}\in\Omega_{v}}p(x_{i}|\Omega_{v})\log p(x_{i}|\Omega_{v})$ (5)
@@
$H_{v}^{A}(X):=\sum\nolimits_{i:\omega_{i}\in\Omega_{v}^{A}}p(x_{i}|\Omega_{v}^{A})\log p(x_{i}|\Omega_{v}^{A})$ (6)
@@
$R_{u}(v)=\widetilde{IG}_{v}(X):=(-p_{v}^{A}\log p_{v}^{A}-p_{v}^{N}\log p_{v}^{N})/(1+\lambda^{-1}|p_{v}^{A}-p_{v}^{N}|)$ (10)
@@ 2.5 Question Selection Via Reward Propagation
$R_{e}(v):=$$\begin{cases}R_{a}(v)&\text{if $v$ is a leaf; otherwise:}\\ p_{v}^{A}R_{e}(v^{A})+p_{v}^{N}R_{e}(v^{N})&\text{if $v$ is an Answerer Node}\\ \frac{1}{m}\sum_{w\in\mathsf{Children}(v)}^{m}R_{e}(w)&\text{if $v$ is a Questioner Node}\\ \end{cases}$
@@ 2.6 UoT Summary
UoT first generates candidate questions $q_{i}^{1},q_{i}^{2},\dots,q_{i}^{m}$ based on the history $h_{i}$ and current possibility set $\Omega_{i}$. Then, we conduct multistep simulation to generate a tree for each candidate question $q_{i}^{n}$. Next, we compute the uncertainty-based rewards $R_{u}(v)$, and propagate over the trees to compute accumulated reward $R_{a}(v)$ and expected reward $R_{e}(v)$. Lastly, the optimal question $q_{i}^{n}$ with highest expected reward will be selected as $q_{i}$ to interact with the Answerer. UoT generates candidate questions $q_{i}^{1},q_{i}^{2},\dots,q_{i}^{m}$ based on history and the current possibility set $\Omega_{i}$. It simulates a tree for each question, calculates uncertainty-based rewards $R_{u}(v)$, and computes expected rewards $R_{e}(v)$. The question $q_{i}^{n}$ with the highest expected reward is chosen for interaction.
@@ 2.7 Extensions and Discussion
Open Set UoT. Recall that in the closed set scenario, the Questioner starts with knowledge of the possibility space $\Omega$. In practice, the possibility space is often unknown, resulting in the open set setting. To adapt UoT to this case, we prompt Questioner to initialize the possibility space $\Omega$ and then reinitialize the possibility set $\Omega_{i}$ according to current history $h_{i}$. Then, the rest of UoT is unchanged. The generalization in open-end answers. The UoT framework enables LLMs to update possibilities after each interaction, including affirmative/negative or open-ended responses. Thus, it can be applied to open-ended answers scenarios. Pruned UoT. To enhance efficiency during simulation, pruning akin to Beam Search can be employed when constructing the simulation trees, which limits the number of paths to explore over the tree to a predetermined size.
@@ 3.1 Experimental Setup
We test various LLMs to evaluate the generality of our method, including Llama-3-70B-Instruct AI@Meta (2024), Mistral-Large Mistral.AI (2024), Gemini-1.5-Pro Reid et al. (2024), Claude-3-Opus Anthropic (2024) and GPT-4 OpenAI (2023b). We also validate the performance of earlier released LLMs (Refer to Appendix D) including Llama 2-70B-Chat Touvron et al. (2023), Cohere Cohere (2023), PaLM 2 Anil et al. (2023), Claude 2 Anthropic (2023) and GPT-3.5-turbo OpenAI (2023a).
Baselines
Direct Prompting (DP) prompts an LLM directly to generate the next response. Planning Prompting (PP) is motivated by Wang et al.Wang et al. (2023). We leverage another LLM to plan the future and, consequently, determine the question to ask. Chain-of-Thought (CoT) Wei et al. (2022) improves reasoning in LLMs by detailing reasoning steps. CoT-SC (Self-Consistency) Wang et al. (2022) an is an ensemble method, explores multiple reasoning paths. We standardize sampling counts for fair computational cost comparison with other methods. Reflexion Shinn et al. (2023) lets agents propose actions and self-assess to foster new ideas. Tree-of-Thoughts (ToT) Yao et al. (2023) enables LLMs to make decisions by exploring and evaluating multiple reasoning paths over a tree structure. We examine ToT under two setups: Original-ToT, which uses the standard approach of generating and evaluating questions, and Adapted-ToT (Ad.-ToT), where we integrate heuristic experience into prompt for question generation and evaluation, focusing on questions that halve the search space. We matched the tree depth to the simulation steps in our UoT method for a fair comparison. We evaluate methods and LLMs in both open set (OS) and closed set (CS) settings. In open set, models are tested without prior knowledge of outcomes; in closed set, they are given complete information about all possible outcomes. For details, see Appendix I.1 for experimental settings and Appendix L for prompts.
@@
Scenarios and Datasets
20 Questions is a game where the answerer thinks of an item and the questioner asks up to 20 yes-or-no questions to guess it. We use two datasets, Common (collected by us, refer to Appendix 6 for more details) and Things Hebart et al. (2019), including 111 and 1854 items separately. In this scenario, the maximal turns is set to 20. In Medical Diagnosis, the doctor needs to ask questions to patients about their symptoms, to determine an accurate diagnosis. We use two datasets: DX Xu et al. (2019), with 104 doctor-patient dialogues and 5 diseases in test set, and MedDG Liu et al. (2022) with over 17K conversations across 15 disease types. We manually selected 500 high-quality samples for evaluation (see Appendix I.3 for selection process). Importantly, Open-ended responses from patient are allowed in MedDG to validate UoT’s generalization in open-ended scenarios. Both datasets are limited to 5 turns. Troubleshooting is a scenario where a customer support technician interacts with customers to identify and resolve faults or issues within computer systems, electronic devices, machinery, or other complex systems. Raghu et al.Raghu et al. (2021) introduce FloDial with 894 dialogues, containing 153 faults and we also conduct the data preprocessing of FloDial (See Appendix I.4 for details). We evaluate using a maximum of 20 turns. The answerer, simulated by GPT-4, is prompted with the patient’s actual disease and conversation details for each case. For more details, refer to Appendix I.2 and see examples of these scenarios in Appendix K.
UoT (Open Set) Setup
We iteratively update LLMs’ perceived possibilities based on conversational history, rather than defining them all upfront. In medical diagnosis and troubleshooting, initial descriptions from symptoms or issues help set up initial possibilities. In the 20-question game, we start with broad inquiries using the Direct Prompting method for the first three rounds to gather more information. The ToT tree structure method employs a similar strategy. Setup details in Appendix I.5.
Evaluation Metrics
To measure efficacy and efficiency, we use: Success Rate (%): $\textbf{SR}=S/T$, where $S$ is the number of successful cases, and $T$ is the total number of cases; Mean Conversation Length in Successful Cases: $\textbf{MSC}=R_{s}/S$, where $R_{s}$ is the total rounds in successful cases; Mean Conversation Length: $\textbf{MCL}=R/T$, where $R$ is the total rounds in all cases. MCL measures efficiency based on the resources used in both successes and failures.
@@ 3.2 Performance
20 Questions
As illustrated in Table 5, for all types of LLMs, those equipped with UoT outperform the baselines in both open set and close settings. Among the methods used on GPT-4 to enhance planning and reasoning, CoT (CS) and PP (CS) show inferior performance even compared to GPT-4 alone. UoT (OS) demonstrates superior performance, with with an average 8.7% improvement than Adapted-ToT (OS) in success rate. Moreover, UoT (CS) achieves the highest success rate, surpassing the second-best Reflexion by an average of 4.3%.
Medical Diagnosis
UoT (CS) outperforms baselines in simplified medical diagnostics, achieving a 97.0% success rate on the DX dataset with GPT-4. On the MedDG dataset, UoT (CS) on Gemini-1.5-Pro and GPT-4 achieve success rates of 81.4% and 88.0%. It also reduces conversation lengths to an average MSC of 2.0 on GPT-4 for DX, lower than 3.5 and 3.0 for DP methods. These results demonstrate the versatility of our UoT in handling both binary and open-ended interactions effectively.
Troubleshooting
UoT (CS) with GPT-4 similarly achieves the highest SR of 67.3%, and the lowest MSC of 7.8. It also shows a remarkable improvement from 43.7% to 67.3% in Success Rate.
Overall Performance
On average, UoT enhances the success rate by 38.1% compared to DP across 5 datasets and 5 different LLMs, including open source and commercial models. Notably, Success Rate increases 46.6% for Llama3-70B. Furthermore, UoT outperforms CoT-SC by 33.8% and Reflexion by 29.9%. Even compared to tree structure methods like Original-ToT and Adapted-ToT, UoT still shows superior performance with gains of 28.3% and 12.4% respectively. Additionally, Pruned UoT, our pruning method to improve efficiency, outperforms Adapted-ToT by 7.36%. Additionally, our study shows that UoT’s one-step planning is effective due to effective reward design and question selection. We limit simulations to three steps for budgetary reasons, balancing efficiency and effectiveness (see Appendix E for further details on simulation depth). To determine whether the differences in success rates between the two methods were statistically significant, we performed a t-test. The results and details are in Appendix H.
@@ 20 Questions
Case Studies and Reliability of GPT-4 as answerer
Figure 3 shows UoT, compared to direct prompting, more effectively reduce uncertainty and narrow down candidates, avoiding overly specific queries. After gaining initial information (e.g., stomach pain), it generates targeted questions about related issues rather than general inquiries. Additionally, GPT-4’s accuracy as answerer is evaluated by analyzing 10% of interactions from each dataset, consistently showing reliable responses. For quantitative details, see Appendix F.
@@ 4 Related Work
Planning and Reasoning of LLMs
LLMs show prowess in planning and reasoning. Wei et al.Wei et al. (2022) introduced CoT prompting for intermediate reasoning; Yao et al.Yao et al. (2023) proposed ToT prompting using DFS/BFS. Besta et al.Besta et al. (2023) present GoT to solve elaborate problems. Feng et al.Feng et al. (2023) illustrated TS-LLM’s tree-search guided decoding. ReAct Yao et al. (2022) offers acting-based prompting, while Reflexion Shinn et al. (2023) enhances this with feedback reflection. Zhou et al.Zhou et al. (2023) unify reasoning and planning.
Decision-making and Information-seeking by LLMs
LLMs have evolved as decision-making tools, with models like LLM+P Liu et al. (2023) and LLM-DP Dagan et al. (2023) combining external planners and LLMs for natural language-based programming. RAP Hao et al. (2023) goes beyond structured language, using LLMs with Monte Carlo Tree Search (MCTS) Chaslot et al. (2008) for dynamic decision-making. This approach is also seen in the work of Zhao et al.Zhao et al. (2023), applying MCTS and LLM knowledge for complex tasks like robot control. However, MCTS struggles in uncertain scenarios due to its reliance on terminal states and specific modules for rewards and action selection. Additionally, to enhance LLMs’ questioning abilities, Deng et al.Deng et al. (2023) introduce the Rephrase and Respond method. AVIS Hu et al. (2023) represents an autonomous visual question answering system that uses external tools. Pan et al.Pan et al. (2023) introduce KwaiAgents for processing queries, following guidelines, and accessing external documents. Frameworks such as MEDIQ Li et al. (2024) and MDAgents Kim et al. (2024) improve the reliability of LLMs in clinical settings by strengthening information-seeking capabilities and agent systems, thereby supporting more realistic diagnostic processes. Bertolazzi et al. Bertolazzi et al. (2023) explore Chatgpt’s information seeking strategy in 20-questions game.
5 Limitation and Future Work